### v1

# Order-agnostic Identifier for Large Language Model-based Generative Recommendation

Abstract.
Leveraging Large Language Models (LLMs) for generative recommendation has attracted significant research interest, where item tokenization is a critical step. It involves assigning item identifiers for LLMs to encode user history and generate the next item. Existing approaches leverage either token-sequence identifiers, representing items as discrete token sequences, or single-token identifiers, using ID or semantic embeddings. Token-sequence identifiers face issues such as the local optima problem in beam search and low generation efficiency due to step-by-step generation. In contrast, single-token identifiers fail to capture rich semantics or encode Collaborative Filtering (CF) information, resulting in suboptimal performance.
To address these issues, we propose two fundamental principles for item identifier design: 1) integrating both CF and semantic information to fully capture multi-dimensional item information, and 2) designing order-agnostic identifiers without token dependency, mitigating the local optima issue and achieving simultaneous generation for generation efficiency. Accordingly, we introduce a novel set identifier paradigm for LLM-based generative recommendation, representing each item as a set of order-agnostic tokens. To implement this paradigm, we propose SETRec, which leverages CF and semantic tokenizers to obtain order-agnostic multi-dimensional tokens. To eliminate token dependency, SETRec uses a sparse attention mask for user history encoding and a query-guided generation mechanism for simultaneous token generation. We instantiate SETRec on T5 and Qwen (from 1.5B to 7B). Extensive experiments on four datasets demonstrate its effectiveness across various scenarios (e.g., full ranking, warm- and cold-start ranking, and various item popularity groups). Moreover, results validate SETRec’s superior efficiency and show promising scalability on cold-start items as model sizes increase.
1. Introduction
Large Language Models (LLMs) have recently demonstrated significant success in personalized recommendation, attracting widespread research interests (Shi et al., 2024; Liao et al., 2024). Surpassing the traditional recommender models, LLMs excel in understanding complex user behaviors and diverse item characteristics due to their rich world knowledge and strong reasoning ability (Touvron et al., 2023). Typically, LLM-based recommenders transform the user’s historical interactions into a token sequence to generate the target item as a recommendation. As shown in Figure 2, a fundamental step of this process is item tokenization, which assigns each item an identifier to enable user history encoding and item generation. Therefore, item tokenization is essential in advancing LLM-based generative recommendation.
Existing item identifiers for LLM-based generative recommendation can be broadly categorized into two groups:
Token-sequence identifier utilizes a discrete token sequence to represent multi-dimensional item information. To generate items, LLMs use beam search to generate the top- item identifiers. Despite the effectiveness, token-sequence identifiers suffer from the 1) local optima issue (Zhou and Hansen, 2005) in the beam search. As illustrated in Figure 1, beam search greedily selects the sequence with top- probabilities at each generation step. However, the initial tokens of the target identifier might not necessarily align with the user preference. As such, the prefix of the target identifier has a low probability and will be pruned by beam search, causing inaccurate results. 2) Low generation efficiency in the autoregressive generation, which requires multiple serial LLM calls, thereby causing unaffordable computing burdens (Lin et al., 2024e) and severely hindering real-world deployments.
Single-token identifier represents each item with a continuous token, i.e., ID embedding or semantic embedding (Liao et al., 2024; Wang et al., 2024d). To recommend items, LLMs first generate the next item embedding, which is then grounded to the item IDs with a linear projection layer, as exemplified by E4SRec (Li et al., 2023) and LITE-LLM4Rec (Wang et al., 2024d). However, using single embedding often yields suboptimal performance. Precisely, ID embeddings rely on sufficient interactions to capture Collaborative Filtering (CF) information, thus being vulnerable to long-tailed users/items. Conversely, semantic embedding overlooks the modeling of CF information that is essential for personalized recommendations.
Facing the above issues, a fundamental question arises: How can we design item identifiers to ensure effective and efficient LLM-based recommendations? Based on the above insights, we posit two principles. 1) Integration of semantic and CF information. Semantic information can harness rich knowledge in LLMs to strengthen the generalization ability (e.g., cold-start recommendation). Meanwhile, CF information leverages user behaviors to enrich the semantic modeling of user preference, enabling effective recommendations for users and items with rich interactions. 2) Order-agnostic Identifier. Representing multi-dimensional item information (e.g., semantic and CF information) with a single token might be suboptimal due to potential conflicts between different dimensions as proven in (Wang et al., 2024a; Zhang et al., 2022) (see empirical evidence in Section 4.3.5). Therefore, it is necessary to utilize a set of tokens to effectively represent items with multi-dimensional information. Nevertheless, multi-dimensional information is not necessarily dependent on each other (e.g., “price” and “category”). Moreover, ordered token sequences can risk the local optima issue. Hence, it is beneficial to disregard token dependencies in identifiers, which further facilitates simultaneous token generation, thus significantly improving inference efficiency.
In this light, we introduce a novel paradigm of set identifier for LLM-based generative recommendation. As shown in Figure 2, it employs a set of order-agnostic tokens to represent each item with CF and semantic information. Nonetheless, it is non-trivial to eliminate token dependencies due to the following challenges:
For user history encoding, the transformed item sequence naturally introduces unnecessary token dependencies (e.g., semantic tokens are dependent on the CF token), which might negatively affect user history encoding.
For the simultaneous generation of order-agnostic identifiers, tokens are independently generated for each dimension (e.g., CF). This necessitates guidance on LLMs to generate tokens aligning well with each information dimension respectively.
Since the tokens for different dimensions are generated independently, the generated set identifier might be invalid items, which requires effective grounding to the existing items.
To this end, we propose SETRec, an effective implementation of the set identifier paradigm. To integrate semantic and CF information, SETRec leverages CF and semantic tokenizers to assign each item with an order-agnostic token set containing CF and semantic embeddings. To eliminate token dependencies, 1) for user history encoding, we propose a special sparse attention mask, which discards the visibility of other tokens within identifiers and retains access to previous identifiers. 2) For simultaneous token generation, we introduce a query-guided generation mechanism, which adopts learnable vectors to guide LLMs to generate the embedding for each specific information dimension. 3) To ground the generated embedding set to existing items, SETRec collects embeddings from all items as grounding heads to obtain the item scores for ranking. We instantiate SETRec on T5 and Qwen and evaluate it on four real-world datasets under various scenarios (e.g., full ranking, warm- and cold-start ranking, and diverse item popularity groups) to demonstrate the effectiveness, efficiency, and generalization ability. Additionally, we evaluate SETRec on Qwen with different model sizes (i.e., 1.5B, 3B, and 7B), exhibiting promising scalability on cold-start items as model size increases.
The main contributions of this work are summarized as follows:
We propose a novel set identifier paradigm for LLM-based generative recommendation, representing each item with a set of order-agnostic tokens integrating semantic and CF information.
We propose SETRec to implement the novel paradigm, which introduces a query-guided generation mechanism with a sparse attention mask to achieve simultaneous generation without token dependencies, significantly boosting inference efficiency.
We instantiate SETRec on T5 and Qwen from 1.5B to 7B. Extensive experiments on four real-world datasets under various settings (e.g., full ranking, warm- and cold-start ranking) validate its effectiveness, efficiency, generalization ability, and scalability.
2. Preliminaries
LLM-based Generative Recommendation. Harnessing LLMs’ strong capabilities, LLM-based generative recommendation aims to use LLMs as recommenders to directly generate personalized recommendations. Formally, given the recommendation data , where is the user’s historical interactions in chronological order and , the target is to utilize a tokenizer to tokenize items into item identifiers , and an LLM-based recommender model to encode the transformed user history and generate next item identifier.
Bridging the language space and the item space, item identifier is a fundamental component for LLMs to encode user history and generate items. Existing identifiers can be divided into two groups:
Token-sequence identifier assigns each item with a discrete token sequence, i.e., , where is the discrete token. Given the user history , it is transformed to an identifier sequence , which is then encoded by LLMs to generate the next identifier via autoregressive generation:
where is the LLM vocabulary. Despite the effectiveness, generating token sequences would result in the local optima issue and inference inefficiency. As shown in Figure 3, continuously increasing the beam size slightly improves recommendation accuracy, but remains inferior to globally optimal results. Worse still, the token-by-token generation requires multiple serial LLM calls, which significantly lowers the inference speed and hinders real-world applications.
Single-token identifier assigns each item with an ID or semantic embedding, i.e., , which is usually obtained by a conventional CF recommender model (e.g., SASRec (Kang and McAuley, 2018)) or a pre-trained semantic extractor (e.g., SentenceT5 (Ni et al., 2021)). Given the transformed user history , it first generates the embedding:
where is the attention layers from the LLM . Based on the generated item embedding , an additional grounding head is added on top of the LLM layers to obtain the scores for all items for ranking. Although it improves inference efficiency by bypassing the token-by-token autoregressive generation, representing items with a single ID embedding struggles with items with fewer interactions while a single semantic embedding overlooks the crucial CF information, thus leading to suboptimal results.
Based on the above insights, we summarize two fundamental principles for identifier designs: 1) integration of semantic and CF information, to leverage rich multi-dimensional item information, and 2) order-agnostic identifier, to eliminate the unnecessary dependencies between tokens associated with an identifier, which can alleviate the local optima issue and improve generation efficiency. In this light, we introduce a novel set identifier paradigm, which employs a set of order-agnostic tokens to represent multi-dimensional item information.
3. SETRec
To implement the set identifier paradigm, we propose a framework called SETRec for effective and efficient LLM-based generative recommendation, including order-agnostic item tokenization and simultaneous item generation as illustrated in Figure 4.
3.1. Order-agnostic Item Tokenization
Meeting the two principles, SETRec leverages a CF and a semantic tokenizer to endow multi-dimensional information into a set of order-agnostic continuous tokens111We do not use discrete tokens in SETRec because discretization inevitably suffers from information loss (Lazebnik and Raginsky, 2008), potentially leading to suboptimal results. as illustrated in Figure 4(b).
CF Tokenizer. As shown in Figure 4(b), we utilize a pre-trained conventional recommender model (e.g., SASRec (Lazebnik and Raginsky, 2008)) with a linear projection layer to obtain item CF embedding , where is the hidden dimension of LLMs. Incorporating CF embeddings encourages LLM-based recommenders to facilitate recommendations for users/items with rich interactions.
Semantic Tokenizer. To fully utilize rich item semantic information, SETRec introduces a semantic tokenizer to obtain a set of semantic embeddings. Specifically, given the item semantic information such as title and categories, we first extract the item semantic representations with a pre-trained semantic extractor (e.g., SentenceT5 (Ni et al., 2021)).
To obtain the semantic embeddings, a straightforward approach is to compress semantic representation into a single latent semantic embedding. Nonetheless, compressing multi-dimensional semantic information (e.g., “brand” and “price”) might suffer from the embedding collapse issue (Guo et al., [n. d.]; Pan et al., 2024), potentially undermining the rich semantic content that distinguishes between items. To prevent this issue, as depicted in Figure 4(b), we tokenize each item into order-agnostic semantic embeddings via an AE:
where denotes the concatenated semantic embeddings representing different latent semantic dimensions, and is the -th semantic embedding. Notably, we utilize a unified AE instead of multiple independent AEs for two considerations: 1) employing a single AE reduces the parameters with an approximate ratio of , which is highly practical; 2) alleviating the training instability that might be caused by multiple encoders’ training (Tang et al., 2023). In addition, to encourage the semantic embeddings to preserve useful information as much as possible, a reconstruction loss is used to train the semantic tokenizer:
where is the reconstructed semantic representation.
Token Corpus. Based on the CF and the semantic tokenizer, we can obtain the set identifier for each item , consisting of a CF embedding and semantic embeddings. We then can collect tokens from all items and obtain the token corpus for each information dimension, i.e., . The collected token corpus is used as the grounding head for effective item grounding (cf. Section 3.2.1).
3.2. Simultaneous Item Generation
To efficiently and effectively generate set identifiers, it is crucial for SETRec to 1) guide LLMs to distinguish different dimensions and generate tokens aligning well with each dimension simultaneously (Section 3.2.1); 2) ground the generated token set to existing items effectively (Section 3.2.1); 3) eliminate the unnecessary dependencies introduced in user history (Section 3.2.2);
3.2.1. Query-guided Generation.
As shown in Figure 4(a), to guide LLMs to generate tokens that align well with the information dimensions, we introduce a set of learnable query vectors , where is the latent dimension of the LLMs, to guide the LLMs to distinguish between information dimensions (e.g., CF and semantic) for token generation. Formally, the generated token for each dimension is obtained via:
where is the learnable query vector to guide LLM generation for the information dimension . Based on Eq. (5), we can collect the generated token for all dimensions and obtain the generated set identifier .
Token Generation Optimization. To achieve accurate item recommendations, we encourage the generated token to align with the target token for every dimension:
where , is the similarity function (e.g., inner product), and is the target item token for the information dimension . Intuitively, Eq. (6) pushes the generated embedding closer to the target embedding and pulls away from other embeddings within the specific information dimension.
Token Generation Grounding. Based on generated tokens obtained via Eq. (5), the next step is to ground them to the existing items. However, this can be challenging since the possible combinations of the tokens from different information dimensions are much larger than the existing item corpus, i.e., . To solve this issue, we introduce a token set grounding strategy, which leverages the token corpus as grounding heads to obtain the item score. Formally, we have
where is adopted from the token corpus . The final item scores are obtained via a linear combination of CF and semantic dimensions, where is a hyper-parameter to balance the strength between CF and semantic dimensions. It is highlighted that the grounding heads for semantic dimensions are extendable to new items, leading to strong generalization ability (cf. Section 4.2).
3.2.2. Sparse Attention Mask.
While simultaneous generation bypasses the sequential generation of item identifier, the flattened user’s historical interactions are still sequentially encoded, inevitably introducing order information of tokens within each identifier (Figure 5(a)). To combat this issue, we introduce a sparse attention mask as illustrated in Figure 5(b). Specifically, for the user’s historical interactions, tokens associated with an identifier are treated as independent from each other (e.g., CF embedding cannot attend to semantic embeddings). However, these tokens can still attend to all tokens in previously interacted items (e.g., a fully attended mask is applied to football when calculating self-attention for tokens in basketball). Therefore, the sparse attention mask ensures the order agnosticism of the set identifier.
Time Complexity Analysis. Moreover, the sparse attention mask can improve the generation efficiency by reducing the duplicate computations of the shared prefix via original attention mask (Figure 5(a)). With information dimensions and historically interacted items, the time complexity for batch generation with the original attention mask is . Remarkably, based on the flattened input with our proposed sparse attention mask, the time complexity reduces to .
3.3. Instantiation
To instantiate SETRec on LLMs, we optimize the CF and semantic tokenizers, learnable query vectors, and LLMs by minimizing:
where is a hyper-parameter to control the strength of the tokenizer training. During inference, SETRec first tokenizes all items into set identifiers and obtain token corpus for each information dimension. Then, to recommend item, SETRec transforms user history into identifier sequence and performs query-guided simultaneous generation with sparse attention mask via Eq. (5) to generate tokens for all information dimensions. Finally, SETRec leverages token corpus as extendable grounding heads to ground the generated token set to the valid items via Eq. (7).
4. Experiment
We carry out extensive experiments on four real-world datasets to answer the following research questions:
RQ1: How does our proposed SETRec perform compared to different identifier baselines on different architectures of LLMs?
RQ2: How do the different components of SETRec (i.e., CF embeddings, semantic embeddings, query vectors, and sparse attention) affect the performance?
RQ3: How does SETRec perform when scaling up the model size and how does SETRec improve the overall performance?
RQ4: How does SETRec perform with different number of semantic embeddings, tokenizer training strength, and semantic strength for inference?
4.1. Experimental Settings
4.1.1. Datasets
We conduct experiments on four real-world datasets across various domains. From Amazon review datasets222https://jmcauley.ucsd.edu/data/amazon/., we adopt three widely used benchmarks 1)Toys, 2) Beauty, and 3) Sports. The three Amazon datasets contain rich user interactions over a specific category of e-commerce products, where each item is associated with rich textual meta information such as title, description, category, and brand. In addition, we use a video games dataset 4) Steam333https://github.com/kang205/SASRec. proposed in (Kang and McAuley, 2018), which contains substantial user interactions on video games with abundant textual semantic information. For all datasets, we follow previous work (Wang et al., 2023) to sort user interactions chronologically according to the timestamps and divide them into training, validation, and testing sets with a ratio of 8:1:1. In addition, we divide the items into warm and cold items444We denote warm- and cold-start items as warm and cold items for brevity., where the items that appear in the training set are warm items, otherwise cold items.
Evaluation. We adopt the widely used metrics Recall@ and NDCG@, where and to evaluate all methods. Additionally, we introduce three different settings that evaluate over 1) all items, 2) warm items only, and 3) cold items only, respectively.
| All | Warm | Cold | Inf. Time (s) | |||||||||||
| Dataset | Method | R@5 | R@10 | N@5 | N@10 | R@5 | R@10 | N@5 | N@10 | R@5 | R@10 | N@5 | N@10 | All Users |
| Toys | DreamRec | 0.0020 | 0.0027 | 0.0015 | 0.0018 | 0.0027 | 0.0039 | 0.0020 | 0.0024 | 0.0066 | 0.0168 | 0.0045 | 0.0082 | 912 |
| E4SRec | 0.0061 | 0.0098 | 0.0051 | 0.0064 | 0.0081 | 0.0128 | 0.0065 | 0.0082 | 0.0065 | 0.0122 | 0.0056 | 0.0078 | 55 | |
| BIGRec | 0.0008 | 0.0013 | 0.0007 | 0.0009 | 0.0014 | 0.0019 | 0.0011 | 0.0013 | 0.0278 | 0.0360 | 0.0196 | 0.0223 | 2,079 | |
[…]
| CID | 0.0044 | 0.0082 | 0.0040 | 0.0053 | 0.0065 | 0.0128 | 0.0049 | 0.0071 | 0.0059 | 0.0111 | 0.0047 | 0.0066 | 810 | |
| SemID | 0.0071 | 0.0108 | 0.0061 | 0.0074 | 0.0086 | 0.0153 | 0.0075 | 0.0100 | 0.0307 | 0.0507 | 0.0220 | 0.0292 | 1,215 | |
[…]
| Beauty | DreamRec | 0.0012 | 0.0025 | 0.0013 | 0.0017 | 0.0016 | 0.0028 | 0.0016 | 0.0019 | 0.0078 | 0.0161 | 0.0065 | 0.0094 | 1,102 |
| E4SRec | 0.0061 | 0.0092 | 0.0052 | 0.0063 | 0.0080 | 0.0121 | 0.0067 | 0.0082 | 0.0072 | 0.0118 | 0.0065 | 0.0077 | 120 | |
| BIGRec | 0.0054 | 0.0064 | 0.0051 | 0.0054 | 0.0008 | 0.0009 | 0.0006 | 0.0008 | 0.0106 | 0.0251 | 0.0095 | 0.0151 | 4,544 | |
[…]
| TIGER | 0.0063 | 0.0098 | 0.0050 | 0.0062 | 0.0086 | 0.0131 | 0.0065 | 0.0082 | 0.0190 | 0.0325 | 0.0130 | 0.0178 | 430 | |
[…]
| Sports | DreamRec | 0.0027 | 0.0044 | 0.0025 | 0.0031 | 0.0032 | 0.0052 | 0.0028 | 0.0035 | 0.0045 | 0.0108 | 0.0026 | 0.0049 | 2,100 |
| E4SRec | 0.0079 | 0.0131 | 0.0075 | 0.0094 | 0.0092 | 0.0154 | 0.0085 | 0.0107 | 0.0031 | 0.0093 | 0.0019 | 0.0039 | 117 | |
| BIGRec | 0.0033 | 0.0042 | 0.0030 | 0.0033 | 0.0001 | 0.0002 | 0.0001 | 0.0001 | 0.0059 | 0.0104 | 0.0043 | 0.0061 | 7,822 | |
| IDGenRec | 0.0087 | 0.0127 | 0.0079 | 0.0092 | 0.0101 | 0.0149 | 0.0091 | 0.0107 | 0.0181 | 0.0302 | 0.0134 | 0.0179 | 1,724 | |
| CID | 0.0077 | 0.0131 | 0.0073 | 0.0092 | 0.0074 | 0.0119 | 0.0045 | 0.0061 | 0.0082 | 0.0149 | 0.0075 | 0.0099 | 2,135 | |
[…]
| TIGER | 0.0085 | 0.0129 | 0.0080 | 0.0095 | 0.0100 | 0.0151 | 0.0091 | 0.0109 | 0.0190 | 0.0310 | 0.0120 | 0.0159 | 481 | |
| LETTER | 0.0077 | 0.0131 | 0.0073 | 0.0092 | 0.0074 | 0.0119 | 0.0045 | 0.0061 | 0.0082 | 0.0149 | 0.0075 | 0.0099 | 481 | |
[…]
| Steam | DreamRec | 0.0029 | 0.0057 | 0.0037 | 0.0046 | 0.0042 | 0.0080 | 0.0045 | 0.0059 | 0.0017 | 0.0029 | 0.0013 | 0.0018 | 4,620 |
| E4SRec | 0.0194 | 0.0351 | 0.0220 | 0.0270 | 0.0312 | 0.0558 | 0.0283 | 0.0370 | 0.0006 | 0.0010 | 0.0006 | 0.0007 | 328 | |
[…]
| IDGenRec | 0.0199 | 0.0307 | 0.0241 | 0.0265 | 0.0309 | 0.0479 | 0.0311 | 0.0363 | 0.0047 | 0.0151 | 0.0039 | 0.0078 | 2,846 | |
[…]
| SemID | 0.0155 | 0.0278 | 0.0192 | 0.0229 | 0.0248 | 0.0443 | 0.0246 | 0.0313 | 0.0017 | 0.0027 | 0.0015 | 0.0018 | 3,605 | |
[…]
4.1.2. Baselines
We compare SETRec with competitive baselines, including single-token identifiers (DreamRec, E4SRec) and token-sequence identifiers (BIGRec, IDGenRec, CID, SemID, TIGER, LETTER). 1) DreamRec (Yang et al., 2024a) is a closely related method that leverages ID embedding to represent each item and adopts a diffusion model to refine the generated ID embedding from LLMs. 2) E4SRec (Li et al., 2023) utilizes a pre-trained CF model to obtain ID embedding, and uses a linear projection layer to obtain the item scores efficiently. 3) BIGRec (Bao et al., 2023) adopts item titles as identifiers, where the tokens are from human vocabulary. 4) IDGenRec (Tan et al., 2024) is a learnable ID generator, which aims to generate concise but informative tags from human vocabulary to represent each item. 5) CID (Hua et al., 2023) leverages hierarchical clustering to obtain token sequence, which utilizes item co-occurrence matrix to obtain identifiers to ensure items with similar interactions share similar tokens. 6) SemID (Hua et al., 2023) also represents items with external token sequence, which is obtained based on the hierarchical item category. 7) TIGER (Rajput et al., 2023) leverages RQ-VAE with codebooks to quantize item semantic information into token sequence with external tokens. The identifier sequentially contains coarse-grained to fine-grained information. 8) LETTER (Wang et al., 2024a) is one of the SOTA item tokenization methods, which incorporates both semantic and CF information into the training of RQ-VAE, achieving identifiers with multi-dimensional information and improved diversity.
4.1.3. Implementation Details
We instantiate all methods on two LLMs with different architectures, i.e., T5-small (Raffel et al., 2020) (encoder-decoder) and Qwen2.5 (Yang et al., 2024b) (decoder-only). Specifically, we adopt Qwen555We denote T5-small and Qwen2.5 as T5 and Qwen for brevity. with different sizes, including 1.5B, 3B, and 7B, for a comprehensive evaluation. To ensure a fair comparison, we set the hidden layer dimensions at 512, 256, and 128 with ReLU activation for methods that adopt AE in tokenizer training, including TIGER, LETTER, and our proposed SETRec. For LLM training, we adopt the same prompt for all methods as “What would the user be likely to purchase next after buying items history?;” for a fair comparison. We fully fine-tune the T5 model and perform parameter-efficient fine-tuning technique LoRA (Hu et al., 2021) for Qwen. All experiments are conducted on four NVIDIA RTX A5000 GPUs. For SETRec, we select , , and from , , and , respectively.
| All | Warm | Cold | Inf. Time(s) | |||||||||||
| Dataset | Method | R@5 | R@10 | N@5 | N@10 | R@5 | R@10 | N@5 | N@10 | R@5 | R@10 | N@5 | N@10 | All Users |
| Toys | DreamRec | 0.0006 | 0.0013 | 0.0005 | 0.0008 | 0.0008 | 0.0019 | 0.0007 | 0.0012 | 0.0076 | 0.0137 | 0.0052 | 0.0074 | 1,093 |
[…]
| BIGRec | 0.0009 | 0.0016 | 0.0009 | 0.0012 | 0.0011 | 0.0013 | 0.0010 | 0.0011 | 0.0194 | 0.0311 | 0.0147 | 0.0191 | 43,304 | |
| IDGenRec | 0.0030 | 0.0053 | 0.0022 | 0.0031 | 0.0043 | 0.0086 | 0.0032 | 0.0048 | 0.0189 | 0.0364 | 0.0161 | 0.0224 | 30,720 | |
| CID | 0.0027 | 0.0047 | 0.0025 | 0.0033 | 0.0055 | 0.0084 | 0.0044 | 0.0056 | 0.0055 | 0.0156 | 0.0044 | 0.0081 | 27,248 | |
| SemID | 0.0024 | 0.0042 | 0.0018 | 0.0024 | 0.0034 | 0.0055 | 0.0026 | 0.0034 | 0.0140 | 0.0275 | 0.0095 | 0.0143 | 32,288 | |
[…]
| LETTER | 0.0057 | 0.0093 | 0.0050 | 0.0064 | 0.0080 | 0.0126 | 0.0066 | 0.0085 | 0.0217 | 0.0416 | 0.0170 | 0.0239 | 13,800 | |
[…]
| Beauty | DreamRec | 0.0007 | 0.0009 | 0.0005 | 0.0005 | 0.0010 | 0.0011 | 0.0007 | 0.0007 | 0.0090 | 0.0167 | 0.0075 | 0.0103 | 1,326 |
[…]
| BIGRec | 0.0006 | 0.0010 | 0.0006 | 0.0007 | 0.0010 | 0.0010 | 0.0008 | 0.0008 | 0.0141 | 0.0246 | 0.0094 | 0.0135 | 29,500 | |
[…]
| CID | 0.0046 | 0.0077 | 0.0040 | 0.0052 | 0.0059 | 0.0107 | 0.0051 | 0.0068 | 0.0075 | 0.0155 | 0.0071 | 0.0096 | 27,792 | |
| SemID | 0.0030 | 0.0045 | 0.0027 | 0.0033 | 0.0050 | 0.0076 | 0.0042 | 0.0052 | 0.0159 | 0.0227 | 0.0116 | 0.0159 | 45,160 | |
| TIGER | 0.0041 | 0.0065 | 0.0032 | 0.0041 | 0.0054 | 0.0085 | 0.0042 | 0.0054 | 0.0083 | 0.0167 | 0.0064 | 0.0091 | 12,600 | |
| LETTER | 0.0040 | 0.0069 | 0.0031 | 0.0042 | 0.0051 | 0.0088 | 0.0039 | 0.0054 | 0.0043 | 0.0129 | 0.0043 | 0.0071 | 12,600 | |
[…]
4.2. Overall Performance (RQ1)
4.2.1. Performance on T5.
The performance comparison between baselines and SETRec instantiated on T5 are shown in Table 1, from which we have the following observations:
Token-sequence identifier (BIGRec, IDGenRec, CID, SemID, TIGER, LETTER) generally performs better than single-token identifier under “all”, “warm”, and “cold” settings. This is reasonable because token-sequence identifier represent each item with multiple tokens, which explicitly encode rich item information into different dimensions.
Among the token-sequence identifiers, methods with external tokens (CID, SemID, TIGER, LETTER) generally outperform those relying on human vocabulary (e.g., BIGRec) under “all” and “warm” settings. This is attributed to their hierarchically structured identifier, where the initial tokens represent coarse-grained semantics while subsequent tokens contain fine-grained semantics. This aligns better with the autoregressive generation process, potentially alleviating the local optima issue (Wang et al., 2024a).
When recommending cold items666The higher values on cold performance are due to the limited number of cold items., methods that merely utilize CF information (DreamRec, E4SRec, and CID) fail to give satisfying results. This is not surprising since CF information depends heavily on substantial interactions for training, thereby struggling with cold items. In contrast, methods that integrate semantics into identifiers (BIGRec, IDGenRec, SemID, TIGER, and LETTER) generalize better on cold-start scenarios (superior performance under “cold” setting). Specifically, BIGRec and IDGenRec tend to have competitive performance. This is reasonable because they utilize readable human vocabulary to represent each item, which better leverages rich world knowledge encoded in LLMs.
SETRec significantly outperform all baselines under “all”, “warm”, and “cold” settings across all four datasets. The superior performance is attributed to 1) the incorporation of both CF and semantic information into a set of tokens, which ensures accurate warm item recommendation and strong generalization on cold items; 2) order agnosticism of identifier, which removes the possibly inaccurate dependencies across different tokens associated with an identifier.
From the perspective of efficiency, SETRec significantly reduces the inference time costs compared to the token-sequence identifiers. SETRec achieves an average 15, 11, 18, and 8 speedup on Toys, Beauty, Sports, and Steam, respectively, compared to token-sequence identifiers. The high efficiency is attributed to the simultaneous generation, which generates multiple tokens at a single LLM call, unlocking the real-world deployment of LLM-based generative recommendation.
4.2.2. Performance on Qwen-1.5B
To evaluate SETRec on decoder-only LLMs, we instantiate SETRec and all baselines on Qwen-1.5B. We present the results on Toys and Beauty777We omit the results with similar observations on other datasets to save space. in Table 2, from which we summarize several key different observations from performance on T5 as follows:
Token-sequence identifiers show limited competitiveness compared to the counterparts on T5. A possible reason is that Qwen-1.5B probably contains richer knowledge within its parameters, which amplifies the knowledge gap between the pre-training and recommendation tasks, thereby hindering its adaptation to recommendation tasks with limited interaction data. Conversely, E4SRec yields competitive performance in most cases. This makes sense because E4SRec removes the original vocabulary head and replaces it with an item projection head, thus facilitating effective adaption to the recommendation tasks.
BIGRec and IDGenRec outperform their T5 counterparts on cold items on Beauty. Because they represent items with human vocabulary, which can leverage the rich world knowledge within Qwen-1.5B for better generalization. On the contrary, identifiers with external tokens have inferior cold performance compared to their T5 counterparts. This is also reasonable since it requires extensive interaction data to train external tokens. Otherwise, it is difficult for it to generalize to cold items accurately due to the low generation probability of these external tokens.
SETRec constantly outperforms baselines, which is consistent with the observations on T5. Notably, SETRec instantiated on Qwen-1.5B steadily surpasses SETRec on T5, especially under the “cold” setting. This validates the strong generalization ability of SETRec on different architectures of LLMs. Moreover, as the LLM size increases, the efficiency improvements over the token-sequence identifiers are more significant, resulting in an average of 20 speedup across the two datasets.
4.3. In-depth Analysis
4.3.1. Ablation Study (RQ2)
To study the effectiveness of each component of SETRec, we separately remove semantic tokens (“w/o Sem”), CF token (“w/o CF”). In addition, we replace learnable query vectors with random frozen vectors (“w/o Query”) and use the original attention mask (“w/o SA”), to evaluate the effect of query vectors and the sparse attention mask, respectively. The results of different ablation variants on T5 and Qwen-1.5B on Toys are presented in Figure 6 and we omit the results on other datasets with similar observations to save space.
From the figures, we can find similar observations on T5 and Qwen that 1) removing each component causes performance drops under “all”, “warm”, and “cold” settings, which validates the effectiveness of each component of SETRec. 2) Discarding semantic tokens drastically degrades the recommendation accuracy under “cold” settings. This demonstrates the necessity of incorporating semantics into identifiers. Interestingly, 3) removing semantic tokens leads to worse performance compared to removing CF token. The possible reason for this is the utilization of multiple semantic tokens to represent each item, which highlights the significance of leveraging multi-dimensional semantic information. This observation is also consistent with the results in (Lin et al., 2024c). Nonetheless, 4) while removing CF tokens for T5 leads to inferior performance on cold items, using CF tokens for Qwen might negatively impact on cold items. A possible reason is that the larger-size Qwen is better at understanding semantics due to its stronger knowledge base encoded in the parameters, making the contribution of CF less significant.
4.3.2. Item Group Analysis (RQ3)
To understand how SETRec improves performance, we evaluate it over items with different popularity. We divide the items into 5 groups according to their frequencies and test the models over each group respectively. The performance comparison between SETRec and two competitive baselines from token-sequence identifiers (LETTER) and single-token identifiers (E4SRec) are reported in Figure 7. We can observe that 1) the performance gradually drops from G1 to G5. This makes sense since the less popular items have fewer interactions for LLMs to learn, thus leading to worse generation probabilities. Besides, 2) E4SRec outperforms LETTER on most popular items (G1) but usually yields inferior performance on unpopular items (G2-G5). This is due to that E4SRec only uses CF information, which relies on substantial interactions and therefore struggle on unpopular items. In contrast, LETTER additionally incorporates semantics into identifiers, thus achieving better generalization on sparse items. 3) SETRec consistently excels both E4SRec and LETTER over all groups. Notably, the improvements over sparse items are more significant, which partially explains the superiority of SETRec regarding overall performance.
4.3.3. Scalability on Model Parameters (RQ3)
To investigate whether SETRec can bring continuous performance when expanding the model parameters, we test SETRec on Qwen with different model sizes (1.5B, 3B, and 7B). Performance comparisons between SETRec, E4SRec, and LETTER on Toys are shown in Table 3. From the results, we can find that 1) SETRec clearly shows continued improvements over cold-start items when the model size scales from 1.5B to 7B, demonstrating promising scalability on cold items. We attribute this to the continued improvements of better semantic understanding by expanding the model parameters. Nonetheless, 2) the performance on the warm items fails to continuously improve, indicating a relatively limited scalability over warm items. This shows that the larger models do not necessarily lead to better CF information understanding, which can also be indicated by the limited improvements of E4SRec under “warm” setting. Besides, 3) LETTER shows weak scalability over the three settings. This is mainly due to the utilization of external tokens, which do not necessarily align with the pre-trained knowledge in LLMs, thus showing limited improvements by expanding the model parameters.
| All | Warm | Cold | |||||
| R@10 | N@10 | R@10 | N@10 | R@10 | N@10 | ||
| 1.5B | LETTER | 0.0093 | 0.0064 | 0.0126 | 0.0085 | 0.0416 | 0.0239 |
| E4SRec | 0.0108 | 0.0072 | 0.0144 | 0.0096 | 0.0235 | 0.0111 | |
| SETRec | 0.0188 | 0.0120 | 0.0236 | 0.0151 | 0.0883 | 0.0507 | |
| 3B | LETTER | 0.0109 | 0.0072 | 0.0151 | 0.0097 | 0.0471 | 0.0236 |
| E4SRec | 0.0096 | 0.0061 | 0.0129 | 0.0081 | 0.0218 | 0.0103 | |
| SETRec | 0.0195 | 0.0123 | 0.0258 | 0.0159 | 0.0964 | 0.0571 | |
| 7B | LETTER | 0.0099 | 0.0061 | 0.0137 | 0.0081 | 0.0406 | 0.0216 |
| E4SRec | 0.0088 | 0.0057 | 0.0114 | 0.0072 | 0.0133 | 0.0065 | |
| SETRec | 0.0194 | 0.0115 | 0.0239 | 0.0140 | 0.1016 | 0.0613 | |
4.3.4. Effect of Semantic Strength (RQ4)
To investigate how semantic information contributes to the performance during inference, we vary from to , where indicates that only CF score is used for ranking, and ranks items based solely on semantic scores (Eq. (7)). From the results reported in Figure 9. we can find that 1) Incorporating semantic information during inference is necessary (inferior performance of than , which facilitates global ranking over multi-dimensional information and lead to strong generalization ability. Notably, 2) incorporating semantic scores brings more significant improvements on cold items, underscoring the critical role of semantic information for zero-shot scenarios. Moreover, 3) Gradually increase to rely solely on semantics (), SETRec maintains competitive performance on warm items, which is probably attributed to the implicit alignment between CF and semantic tokens during training.
4.3.5. Hyper-parameter Sensitivity (RQ4)
We further study the hyper-parameter sensitivity to facilitate SETRec application.
Effect of . We vary the strength of AE loss for SETRec training and present the results on Toys in Figure 8(a-b). We can observe that 1) the performance is overall improved when is increased from to , which validates the effectiveness of reconstruction loss that encourages AE to preserve useful information in the latent space. Nonetheless, 2) while continuously increasing generally gives better performance on cold-start items, it might hurt the performance under “warm” setting. Based on the empirical results, we recommend setting ranging from to .
Effect of . We change the number of semantic tokens from to to investigate how affects the performance. From the results shown in Figure 8(c-d), we can find that 1) gradually increasing semantic tokens generally improves the performance, which validates the effectiveness of incorporating multiple tokens to mitigate the potential information conflicts (Wang et al., 2024a) and embedding collapse issue (Guo et al., [n. d.]). However, 2) blindly increasing the number of semantic tokens might hurt the performance (decreased performance from to ). This is reasonable since it is non-trivial to recover the category-level preference aligning well with the real-world scenarios. Similar observations are also seen in (Lin et al., 2024b) and (Lin et al., 2024d).
5. Related Work
LLM-based Recommendation. Harnessing LLMs for recommendations has garnered substantial attention across both academia and industry (Sun et al., 2024; Zhang et al., 2024c; Fu et al., 2024; Zhang et al., 2024c; Shi et al., 2024; Zhao et al., 2024; Zheng et al., 2024a; Li et al., 2024). Existing studies on LLM-based recommendations can be grouped into two research lines. 1) LLMs for discriminative recommendation, which typically aims to leverage LLMs to assist the conventional recommender models (Wang et al., 2024c; Ren et al., 2024a; Chen et al., 2024b; Wu et al., 2024). Harnessing LLMs’ strong reasoning ability, this line of work usually involves LLMs in different steps of recommendation pipelines (Lin et al., 2024a; Zhang et al., 2024a, b; Wang et al., 2024b) such as feature engineering (Xi et al., 2024; Ren et al., 2024b; Xi et al., 2024) and feature encoder (Chen et al., 2024a). 2) LLMs for generative recommendation, which regards LLMs as recommenders to directly generate items (Liao et al., 2024; Kim et al., 2024; Lin et al., 2024c; Wang et al., 2024f; Li et al., 2024). To build LLMs for generative recommendation, a key step is item tokenization, where each item is assigned an identifier for LLMs to encode user history and generate the next item. In this work, we critically analyze the fundamental principles of identifier design to achieve effective and efficient LLM-based generative recommendation.
Identifier for LLM-based Recommendation. Existing identifier designs can be broadly categorized into two types: 1) token-sequence identifiers represent each item with a discrete token sequence. Under this group of work, prior effort has been made to utilize tokens in human vocabulary (i.e., tokens that are included in the LLM vocabulary). Previous work leverages items’ textual information such as titles (Bao et al., 2023), descriptions (Cui et al., 2022), and tags (Tan et al., 2024), aiming to utilize knowledge encoded in LLMs. More recently, utilizing external tokens for identifiers has attracted extensive attention due to its potential to include hierarchical information (Zhu et al., 2024; Zheng et al., 2024b; Wang et al., 2024e). To achieve this, existing work usually adopts hierarchical clustering or RQ-VAE (Lee et al., 2022) to obtain tokens in different granularity. Despite the effectiveness, token-sequence identifiers suffer from local optima issue and inference inefficiency. To improve efficiency, 2) single-token identifiers are proposed to represent each item with ID or semantic embedding (Li et al., 2023; Wang et al., 2024d). Nonetheless, existing work neither has poor generalization ability nor fails to capture CF information, thus leading to suboptimal results. In this work, we propose a novel set identifier paradigm, which employs a set of order-agnostic CF and semantic tokens. Two concurrent studies explore set identifiers for generative retrieval (Zeng et al., 2024; Zhang et al., 2024b), yet they still preserve token dependencies and heavily rely on autoregressive generation. Differently, our proposed paradigm achieves simultaneous generation without token dependencies, significantly enhancing generation efficiency.
6. Conclusion
In this work, we revealed inherent issues of existing identifiers for LLM-based generative recommendation, i.e., inadequate information, local optima, and generation inefficiency. We then summarized two principles for identifier design, i.e., 1) integration of both CF and semantic information, and 2) order-agnostic identifiers. Meeting the two principles, we introduced a novel set identifier paradigm, which employs order-agnostic set identifiers to encode user history and generate the set identifier simultaneously. To implement this paradigm, we proposed SETRec, which uses CF and semantic tokenizers to obtain a set of CF and semantic tokens. To remove token dependencies, we introduced a sparse attention mask for user history encoding and a query-guided generation mechanism for simultaneous generation. Empirical results on four datasets across various scenarios demonstrated the effectiveness, efficiency, generalization ability, and scalability of SETRec.
This work underscores the order agnosticism and multi-dimensional information utilization for identifier design, paving the way for numerous promising avenues for future research. 1) To better align with the pre-training tasks and fully utilize the knowledge within LLMs, it is worth exploring how discrete set identifiers (i.e., a set of order-agnostic discrete tokens) perform on generative recommendation. 2) While SETRec shows strong generalization ability in challenging scenarios such as unpopular item groups, it is worthwhile to apply SETRec for open-ended recommendation with open-domain user behaviors.